# IterNorm: Fast Iterative Normalization

### v1

@@ Abstract
Transformer-based large language models are a memory-bound model whose operation is based on a large amount of data that are marginally reused. Thus, the data movement between a host and accelerator likely dictates the total wall-clock time. Layer normalization is one of the key workloads in the transformer model, following each of multi-head attention and feed-forward network blocks. To reduce data movement, layer normalization needs to be performed on the same chip as the matrix-matrix multiplication engine. To this end, we introduce an iterative L2-normalization method for 1D input (IterNorm), ensuring fast convergence to the steady-state solution within five iteration steps and high precision, outperforming the fast inverse square root algorithm in six out of nine cases for FP32 and five out of nine for BFloat16 across the embedding lengths used in the OPT models. Implemented in 32/28nm CMOS, the IterNorm macro normalizes $d$-dimensional vectors, where $64\leq d\leq 1024$, with a latency of 112-227 cycles at 100MHz/1.05V.
@@
IterNorm, layer normalization, fast convergence, large language models
@@ I Introduction
LLMs based on transformer are memory-bound models that depend on a large amount of data but very limited reuse of them for operations (cf. convolutional neural networks as representative compute-bound models) [5]. The large amount of data needed and their limited reuse cause significant data traffic between a main memory (DRAM) and processor, so that the overall operational wall-clock time is dictated by the memory bandwidth rather than the processor performance. Thus, graphics processing units (GPUs) equipped with high bandwidth memories are widely used to accelerate LLMs. As an alternative to GPUs, various processors have recently been proposed to mainly accelerate MatMul operations and/or activation functions at a lower power than GPUs, including computing-in-memory units such as Function-In-Memory DRAM (FIMDRAM) [6], and GDDR6-based accelerator-in-memory (AiM) [7]. However, most of them rely on their host for layer normalization. That is, the output from each sub-block should be sent to the main memory for layer normalization, which leads to data traffic, and thus significant latency and power consumption.
@@
The layer normalization algorithm for on-chip implementation needs to avoid vanilla division and square root and to be generic to apply to various floating-point (FP) formats with low power/area overheads and operational latency. To this end, we propose a fast iterative normalization algorithm (IterNorm) that is based on a high-dimensional dynamic system with a few fixed point. One of them represents the L2-normalized 1D vector in the hyper-space, which can be attained by appropriately setting the initial point in the hyper-space. IterNorm is a division and square root operation-free L2-normalization algorithm, rendering it suitable for power- and area-efficient on-chip implementations. It is based on a sold theoretical ground and applicable to various FP formats unlike previous methods tailored to specific FP formats [12]. It highlights fast convergence toward the fixed point (L2-normalized vector) in the hyperspace within five iteration steps, and thus low operational latency.
@@
  • We also share our digital implementation of the IterNorm macro with detailed explanation and its power and area overheads in 32/28nm CMOS technology.
The rest of this paper is organized as follows. Sec. II explains the key theories of dynamic systems that prototype IterNorm. Sec. III elaborates IterNorm and the system initialization that ensures fast convergence. Sec. IV proposes the IterNorm macro architecture. Sec. V evaluates IterNorm and the macro, and Sec. VI compares our work with previous ones. Sec. VII concludes our work.
@@ III-A Iterative normalization method
which asymptotically converges towards $a_{\infty}=1/\left\|\boldsymbol{y}\right\|_{2}$ with a positive $a_{0}$ and sufficiently small $\lambda$. We refer to this L2-normalization method with the replacement of Step 2 by this iterative normalization as IterNorm. The pseudocode for IterNorm-based layer normalization is shown in Algorithm 1.
@@
Initialization: $\tilde{\boldsymbol{y}}\leftarrow\tilde{\boldsymbol{y}}_{0}$; $\Delta a\leftarrow\delta_{0}\left(>\delta_{\textrm{max}}\right)$; $a\leftarrow a_{0}$
@@
Algorithm 1 IterNorm-based layer normalization.
Figure 1: (a) Architecture of the IterNorm macro. (b) Data organization in the Input buffer. (c) Block diagram of the Add block equipped with total nine 8-input adder trees.
@@ III-B Initialization and update-rate setting
The more the iteration steps for IterNorm, the larger the wall-clock time. To shorten the iteration, we should use the initial value $a_{0}$ close to $a_{\infty}$$(=1/\left\|\boldsymbol{y}\right\|_{2})$ and update-rate $\lambda$ such that (i) $\lambda$ is sufficiently large for $a_{\infty}$ to be attained with the minimum iteration steps (fast convergence) but (ii) sufficiently small to avoid an intolerable error in approximation.
Initialization of $\boldsymbol{a}$: Let $m$ be $\left\|\boldsymbol{y}\right\|_{2}^{2}$, which is once evaluated for IterNorm as shown in Algorithm 1. We initialize $a$ using the exponent of $m$, $E\left(m\right)$, as follows.
@@
The convergence rate is determined by the exponent on the right-hand side of Eq. (9). For fast convergence, the exponential term should fall below a tolerable error value $\delta_{c}\sim 0$ within a few iteration steps $n_{c}$, leading to the following inequality, $\lambda>-\left(2mn_{c}\right)^{-1}\ln\delta_{c}$. We set $\delta_{c}$ and $n_{c}$ to $10^{-3}$ and $5$, respectively, so that we have $\lambda>0.69m^{-1}$. As such, the calculation of $m^{-1}$ needs a division operation, which is avoided in IterNorm. Because the exponent of $m$, i.e., $E\left(m\right)$ is known, the range of $m^{-1}$ is readily available, $0.5\cdot 2^{-\left(E\left(m\right)-\text{bias}\right)}<m^{-1}\leq 2^{-\left(E\left(m\right)-\text{bias}\right)}$. Therefore, we approximate the condition of $\lambda$ for $a$ to converge within $n_{c}\left(=5\right)$ iteration steps.
@@
IV IterNorm macro design
@@
The IterNorm macro implements the IterNorm-based layer normalization algorithm for a $d$-long input vector $\boldsymbol{x}=\left[x_{0},x_{1},\cdots,x_{d-1}\right]$ with scale parameters $\boldsymbol{\gamma}=\left[\gamma_{0},\gamma_{1},\cdots,\gamma_{d-1}\right]$ and shift parameters $\boldsymbol{\beta}=\left[\beta_{0},\beta_{1},\cdots,\beta_{d-1}\right]$, which outputs layer-normalized input $\boldsymbol{z}=\left[z_{0},z_{1},\cdots,z_{d-1}\right]$. Fig. 1a shows a block diagram of the IterNorm macro proposed. The Input buffer of eight parallel banks ($n_{b}=8$) buffers a $d$-long input vector, and thus the input length $d$ is limited by the buffer size (Fig. 1b). Because each bank can stores $16\times 8$ input elements ($h_{b}=16$ and $w_{b}=8$), the IterNorm macro can handle $d=1024$ maximally for a single input, i.e., $d_{\text{max}}=1024$. A $d$-long input vector $\boldsymbol{x}$ is buffered over multiple banks such that, in a bank $b$ out of total eight banks ($n_{b}=8$), its row $i$ stores $x[w_{b}(b+n_{b}i):w_{b}(b+n_{b}i+1)-1]$ as illustrated in Fig. 1b. Because eight parallel banks share a read pointer, $x[n_{b}w_{b}i:n_{b}w_{b}(i+1)-1]$ is read at a time. Instead, multiple ($\lfloor d_{\text{max}}/d\rfloor$) input vectors can be buffered and sequentially normalized. Note that, to maintain $d_{\text{max}}=1024$ for FP32/16 and BFloat16, the IterNorm macro for FP32 uses the Input buffer twice as large as that for FP16 and BFloat16. Additionally, the Mul and Add blocks are tailored to each data format by using format-specific multipliers and adders but with the same latency of two clock cycles.
@@ V-A Computational precision and convergence rate
Figure 3: Approximation precision of IterNorm for various input lengths $d$ in (a) FP32, (b) FP16, and (c) BFloat16. The insets show the distribution of errors for $d=384$ over 1,000 input vectors.
We evaluated the performance of the IterNorm macro implemented in a Xilinx Virtex-7 FPGA. We applied the IterNorm-based normalization to random vectors of different lengths $(64\leq d\leq 1024)$ in FP32/FP16/BFloat16. For each length and each data format, we used 1,000 random vectors sampled from a uniform distribution in the range $(-1,1)$ as input vectors. The number of iteration steps was set to 5. We used the absolute deviation of our results from the ground truth (absolute error) as a measure of computational precision. The ground truth values were calculated by applying the layer-normalization function in PyTorch (1.12.1) [13] to the same random vectors using a CPU. Fig. 3 shows the evaluation results. The average (maximum) absolute errors for FP32, FP16, and BFloat16 are $2.23\times 10^{-4}(5.0\times 10^{-1})$, $5.26\times 10^{-4}(4.9\times 10^{-1})$, and $3.07\times 10^{-3}(6.8\times 10^{-1})$, respectively. The maximum error cases marginally occurred as shown in the histograms in the insets of Fig. 3.
LLMs use long embedding vectors, as seen in the OPT models, with the largest (OPT-175B) utilizing 12,288-dimensional embeddings [14]. We further evaluated the precision of IterNorm for the embedding lenghts used in the OPT models ($288768\leq d\leq 12,$ 288) and compared with the layer normalization method based on the fast inverse square root (FISR) algorithm [12]. Since FISR is designed for FP formats with an 8b exponent, we limit our comparison to FP32 and BFloat16. The results are shown in Table I. In FP32, IterNorm outperforms the FISR-based method in terms of average precision in six out of nine cases while, in BFloat16, it does so in five out of nine cases.
Figure 4: Average absolute errors of IterNorm in FP32, FP16, and BFloat16 with the number of iteration steps.
TABLE I: Precision comparison between IterNorm and FISR
@@
To identify the covergence rate, we measured the average absolute error by varying the numbers of iteration steps for $d=1024$ for FP32, FP16, and BFloat16. Fig. 4 plots the evaluation results, where each data point was acquired from 1,000 trials. IterNorm in FP16 and BFloat16 ensures fast convergence within five iteration steps while that in FP32 needs a few additional iteration steps until convergence. Nevertheless, the error after five steps is close to the steady state error and far below the steady state errors for FP16 and BFloat16. Note that, in all formats, these errors after five steps may be sufficiently low to avoid LLM-level performance degradation on some text generation tasks as addressed in Sec. V-D.
@@ V-B Operational latency
Figure 5: Measured latency of IterNorm (five iteration steps) with input length $d$.
We evaluated the IterNorm macro latency by varying input length $d$. The results are plotted in Fig. 5. Note that the latency does not rely on the data format. The latency scales with input length $d$. This is because the major steps addressed in Sec. IV work on a chunk of $n_{b}w_{b}(=64)$ input elements at a time, so that the latency scales with $\lceil d/\left(n_{b}w_{b}\right)\rceil$. They include the mean calculation, mean-shift operation, inner product of $\boldsymbol{y}$ with itself, and scale and shift operations.
V-C IterNorm macro in 32/28nm CMOS
TABLE II: Synthesis results for the IterNorm macros in FP32, FP16, and BFloat16
@@ V-C IterNorm macro in 32/28nm CMOS
Figure 6: Area breakdowns for the IterNorm macro for (a) FP32, (b) FP16, and (c) BFloat16. Power breakdowns for (d) FP32, (e) FP16, and (f) BFloat16.
TABLE III: Comparison between the IterNorm macro and previous implementations of layer normalization
Implementation Method Operations Data format Area Power Clock frequency [8] 65nm CMOS approximate SQRT addition, division, bit shift INT32 68.3 $\textrm{mm}^{2}$ 2.0 W 143 MHz [9] 7nm CMOS approximate 1/SQRT multiplication, addition INT32 FP32 FP16 1008.9 ${\mu}m^{2}$ 1133.6 ${\mu}m^{2}$ 498.4 ${\mu}m^{2}$ 59.1 $\mu$W 43.7 $\mu$W 25.0 $\mu$W - [10] 28nm CMOS FISR multiplication, addition, bit shift BFloat16 - - 1 GHz [11] 28nm CMOS layer normalization w/ dynamic compress multiplication, addition, bit shift INT8 - - 1 GHz Ours 32/28nm CMOS IterNorm multiplication, addition FP32 FP16 BFloat16 2.4 (1.7) $\textrm{mm}^{2}$ 1.1 (0.8) $\textrm{mm}^{2}$ 1.0 (0.8) $\textrm{mm}^{2}$ 22.9 mW 8.4 mW 7.3 mW 100 MHz
@@
We finally synthesized the IterNorm macro for each of FP32/FP16/BFloat16 using the Synopsys SAED 32/28nm technology PDK with a supply voltage of 1.05V and a clock frequency $f_{\text{clk}}$ of 100MHz. We used the Design Compiler V-2023.12-SP5. Note that the configuration of Input buffer banks for all formats follows the generic architecture (use of eight Input buffer banks, each of which stores $16\times 8$ input elements) explained in Sec. IV. The synthesis results are summarized in Table II. As such, the macro for FP32 needs on-chip memory (96.5 kib) twice as large as those (48.3 kib) for FP16 and BFloat16. For FP32, each of the Input, $\gamma$, and $\beta$ buffers uses 32 kib to store 1024 elements maximally, and the partial sum buffer uses 0.5 kib to maximally store 16 partial sums. For FP16 and BFloat16, the memory usage is half that for FP32 such that the Input, $\gamma$, and $\beta$ buffers use 48 kib in total, and the partial sum buffer 0.25 kib.
@@
As shown in Table II, the macro areas for FP32, FP16, and BFloat16 are 2.4, 1.1, and 1.0 $\textrm{mm}^{2}$, respectively. The area breakdown for each format is shown in Figs.6a-c. For all formats, the memory (Input/$\gamma$/$\beta$ and partial sum buffers) occupies the largest area in the macro, which is followed by the logic area including FP multipliers and adders. Although we considered FP multipliers and adders dedicated solely to IterNorm, IterNorm can use them in the MatMul block co-integrated on the same die. Therefore, the actual area of the IterNorm macro likely excludes the multiplier and adder areas, which is also listed in Table II. The operational power is also primarily determined by the FP multipliers and adders, resulting in the highest (lowest) power consumption for FP32 (BFloat16) as identified in Table II and power breakdown for each format in Figs.6d-f.

### v2

# IterL2Norm: Fast Iterative L2-Normalization

@@ Abstract
Transformer-based large language models are a memory-bound model whose operation is based on a large amount of data that are marginally reused. Thus, the data movement between a host and accelerator likely dictates the total wall-clock time. Layer normalization is one of the key workloads in the transformer model, following each of multi-head attention and feed-forward network blocks. To reduce data movement, layer normalization needs to be performed on the same chip as the matrix-matrix multiplication engine. To this end, we introduce an iterative L2-normalization method for 1D input (IterL2Norm), ensuring fast convergence to the steady-state solution within five iteration steps and high precision, outperforming the fast inverse square root algorithm in six out of nine cases for FP32 and five out of nine for BFloat16 across the embedding lengths used in the OPT models. Implemented in 32/28nm CMOS, the IterL2Norm macro normalizes $d$-dimensional vectors, where $64\leq d\leq 1024$, with a latency of 116-227 cycles at 100MHz/1.05V.
@@
IterL2Norm, layer normalization, fast convergence, large language models
@@ I Introduction
LLMs based on transformer are memory-bound models that depend on a large amount of data but very limited reuse of them for operations (cf. convolutional neural networks as representative compute-bound models) [5]. The large amount of data needed and their limited reuse cause significant data traffic between a main memory (DRAM) and processor, so that the overall operational wall-clock time is dictated by the memory bandwidth rather than the processor performance. Thus, graphics processing units (GPUs) equipped with high bandwidth memories are widely used to accelerate LLMs. As an alternative to GPUs, various processors have recently been proposed to mainly accelerate MatMul operations and/or activation functions at a lower power than GPUs, including computing-in-memory units such as Function-In-Memory DRAM (FIMDRAM) [6], and GDDR6-based accelerator-in-memory (AiM) [7]. However, most of them rely on their host for layer normalization. That is, the output from each sub-block should be sent to the main memory for layer normalization, which leads to data traffic, and thus significant latency and power consumption [5].
@@
The layer normalization algorithm for on-chip implementation needs to avoid vanilla division and square root and to be generic to apply to various floating-point (FP) formats with low power/area overheads and operational latency. To this end, we propose a fast iterative L2-normalization algorithm (IterL2Norm) that is based on a high-dimensional dynamic system with a few fixed point. One of them represents the L2-normalized 1D vector in the hyper-space, which can be attained by appropriately setting the initial point in the hyper-space. IterL2Norm is a division and square root operation-free L2-normalization algorithm, rendering it suitable for power- and area-efficient on-chip implementations. It is based on a sold theoretical ground and applicable to various FP formats unlike previous methods tailored to specific FP formats [12]. It highlights fast convergence toward the fixed point (L2-normalized vector) in the hyperspace within five iteration steps, and thus low operational latency.
@@
  • We also share our digital implementation of the IterL2Norm macro with detailed explanation and its power and area overheads in 32/28nm CMOS technology.
The rest of this paper is organized as follows. Sec. II explains the key theories of dynamic systems that prototype IterL2Norm. Sec. III elaborates IterL2Norm and the system initialization that ensures fast convergence. Sec. IV proposes the IterL2Norm macro architecture. Sec. V evaluates IterL2Norm and the macro, and Sec. VI compares our work with previous ones. Sec. VII concludes our work.
@@ III-A Iterative normalization method
which asymptotically converges towards $a_{\infty}=1/\left\|\boldsymbol{y}\right\|_{2}$ with a positive $a_{0}$ and sufficiently small $\lambda$. We refer to this L2-normalization method with the replacement of Step 2 by this iterative normalization as IterL2Norm. The pseudocode for IterL2Norm-based layer normalization is shown in Algorithm 1.
@@
Initialization: $\Delta a\leftarrow\delta_{0}\left(>\delta_{\textrm{max}}\right)$; $a\leftarrow a_{0}$
@@
Algorithm 1 IterL2Norm-based layer normalization.
Figure 1: (a) Architecture of the IterL2Norm macro. (b) Data organization in the Input buffer. (c) Block diagram of the Add block equipped with total nine 8-input adder trees.
@@ III-B Initialization and update-rate setting
The more the iteration steps for IterL2Norm, the larger the wall-clock time. To shorten the iteration, we should use the initial value $a_{0}$ close to $a_{\infty}$$(=1/\left\|\boldsymbol{y}\right\|_{2})$ and update-rate $\lambda$ such that (i) $\lambda$ is sufficiently large for $a_{\infty}$ to be attained with the minimum iteration steps (fast convergence) but (ii) sufficiently small to avoid an intolerable error in approximation.
Initialization of $\boldsymbol{a}$: Let $m$ be $\left\|\boldsymbol{y}\right\|_{2}^{2}$, which is once evaluated for IterL2Norm as shown in Algorithm 1. We initialize $a$ using the exponent of $m$, $E\left(m\right)$, as follows.
@@
The convergence rate is determined by the exponent on the right-hand side of Eq. (9). For fast convergence, the exponential term should fall below a tolerable error value $\delta_{c}\sim 0$ within a few iteration steps $n_{c}$, leading to the following inequality, $\lambda>-\left(2mn_{c}\right)^{-1}\ln\delta_{c}$. We set $\delta_{c}$ and $n_{c}$ to $10^{-3}$ and $5$, respectively, so that we have $\lambda>0.69m^{-1}$. As such, the calculation of $m^{-1}$ needs a division operation, which is avoided in IterL2Norm. Because the exponent of $m$, i.e., $E\left(m\right)$ is known, the range of $m^{-1}$ is readily available, $0.5\cdot 2^{-\left(E\left(m\right)-\text{bias}\right)}<m^{-1}\leq 2^{-\left(E\left(m\right)-\text{bias}\right)}$. Therefore, we approximate the condition of $\lambda$ for $a$ to converge within $n_{c}\left(=5\right)$ iteration steps.
@@
IV IterL2Norm macro design
@@
The IterL2Norm macro implements the IterL2Norm-based layer normalization algorithm for a $d$-long input vector $\boldsymbol{x}=\left[x_{0},x_{1},\cdots,x_{d-1}\right]$ with scale parameters $\boldsymbol{\gamma}=\left[\gamma_{0},\gamma_{1},\cdots,\gamma_{d-1}\right]$ and shift parameters $\boldsymbol{\beta}=\left[\beta_{0},\beta_{1},\cdots,\beta_{d-1}\right]$, which outputs layer-normalized input $\boldsymbol{z}=\left[z_{0},z_{1},\cdots,z_{d-1}\right]$. Fig. 1a shows a block diagram of the IterL2Norm macro proposed. The Input buffer of eight parallel banks ($n_{b}=8$) buffers a $d$-long input vector, and thus the input length $d$ is limited by the buffer size (Fig. 1b). Because each bank can stores $16\times 8$ input elements ($h_{b}=16$ and $w_{b}=8$), the IterL2Norm macro can handle $d=1024$ maximally for a single input, i.e., $d_{\text{max}}=1024$. A $d$-long input vector $\boldsymbol{x}$ is buffered over multiple banks such that, in a bank $b$ out of total eight banks ($n_{b}=8$), its row $i$ stores $x[w_{b}(b+n_{b}i):w_{b}(b+n_{b}i+1)-1]$ as illustrated in Fig. 1b. Because eight parallel banks share a read pointer, $x[n_{b}w_{b}i:n_{b}w_{b}(i+1)-1]$ is read at a time. Instead, multiple ($\lfloor d_{\text{max}}/d\rfloor$) input vectors can be buffered and sequentially normalized. Note that, to maintain $d_{\text{max}}=1024$ for FP32/16 and BFloat16, the IterL2Norm macro for FP32 uses the Input buffer twice as large as that for FP16 and BFloat16. Additionally, the Mul and Add blocks are tailored to each data format by using format-specific multipliers and adders but with the same latency of two clock cycles.
@@ V-A Computational precision and convergence rate
Figure 3: Approximation precision of IterL2Norm for various input lengths $d$ in (a) FP32, (b) FP16, and (c) BFloat16. The insets show the distribution of errors for $d=384$ over 1,000 input vectors.
We evaluated the performance of the IterL2Norm macro implemented in a Xilinx Virtex-7 FPGA. We applied the IterL2Norm-based normalization to random vectors of different lengths $(64\leq d\leq 1024)$ in FP32/FP16/BFloat16. For each length and each data format, we used 1,000 random vectors sampled from a uniform distribution in the range $(-1,1)$ as input vectors. The number of iteration steps was set to 5. We used the absolute deviation of our results from the ground truth (absolute error) as a measure of computational precision. The ground truth values were calculated by applying the layer-normalization function in PyTorch (1.12.1) [13] to the same random vectors using a CPU. Fig. 3 shows the evaluation results. The average (maximum) absolute errors for FP32, FP16, and BFloat16 are $2.23\times 10^{-4}(5.0\times 10^{-1})$, $5.26\times 10^{-4}(4.9\times 10^{-1})$, and $3.07\times 10^{-3}(6.8\times 10^{-1})$, respectively. The maximum error cases marginally occurred as shown in the histograms in the insets of Fig. 3.
LLMs use long embedding vectors, as seen in the OPT models, with the largest (OPT-175B) utilizing 12,288-dimensional embeddings [14]. We further evaluated the precision of IterL2Norm for the embedding lengths used in the OPT models ($288768\leq d\leq 12,$ 288) and compared with the layer normalization method based on the fast inverse square root (FISR) algorithm [12]. Since FISR is designed for FP formats with an 8b exponent, we limit our comparison to FP32 and BFloat16. The results are shown in Table I. In FP32, IterL2Norm outperforms the FISR-based method in terms of average precision in six out of nine cases while, in BFloat16, it does so in five out of nine cases.
Figure 4: Average absolute errors of IterL2Norm in FP32, FP16, and BFloat16 with the number of iteration steps.
TABLE I: Precision comparison between IterL2Norm and FISR
@@
To identify the covergence rate, we measured the average absolute error by varying the numbers of iteration steps for $d=1024$ for FP32, FP16, and BFloat16. Fig. 4 plots the evaluation results, where each data point was acquired from 1,000 trials. IterL2Norm in FP16 and BFloat16 ensures fast convergence within five iteration steps while that in FP32 needs a few additional iteration steps until convergence. Nevertheless, the error after five steps is close to the steady state error and far below the steady state errors for FP16 and BFloat16. Note that, in all formats, these errors after five steps may be sufficiently low to avoid LLM-level performance degradation on some text generation tasks as addressed in Sec. V-D.
@@ V-B Operational latency
Figure 5: Measured latency of IterL2Norm (five iteration steps) with input length $d$.
We evaluated the IterL2Norm macro latency by varying input length $d$. The results are plotted in Fig. 5. Note that the latency does not rely on the data format. The latency scales with the number of chunks $\lceil d/\left(n_{b}w_{b}\right)\rceil$ of the input length $d$. This is because the major steps addressed in Sec. IV work on a chunk of $n_{b}w_{b}(=64)$ input elements at a time, so that the latency scales with $\lceil d/\left(n_{b}w_{b}\right)\rceil$. They include the mean calculation, mean-shift operation, inner product of $\boldsymbol{y}$ with itself, and scale and shift operations.
V-C IterL2Norm macro in 32/28nm CMOS
TABLE II: Synthesis results for the IterL2Norm macros in FP32, FP16, and BFloat16
@@ V-C IterL2Norm macro in 32/28nm CMOS
Figure 6: Area breakdowns for the IterL2Norm macro for (a) FP32, (b) FP16, and (c) BFloat16. Power breakdowns for (d) FP32, (e) FP16, and (f) BFloat16.
TABLE III: Comparison between the IterL2Norm macro and previous implementations of layer normalization
Implementation Method Operations Data format Area Power Clock frequency [8] 65nm CMOS approximate SQRT addition, division, bit shift INT32 68.3 $\textrm{mm}^{2}$ 2.0 W 143 MHz [9] 7nm CMOS approximate 1/SQRT multiplication, addition INT32 FP32 FP16 1008.9 ${\mu}m^{2}$ 1133.6 ${\mu}m^{2}$ 498.4 ${\mu}m^{2}$ 59.1 $\mu$W 43.7 $\mu$W 25.0 $\mu$W - [10] 28nm CMOS FISR multiplication, addition, bit shift BFloat16 - - 1 GHz [11] 28nm CMOS layer normalization w/ dynamic compress multiplication, addition, bit shift INT8 - - 1 GHz Ours 32/28nm CMOS IterL2Norm multiplication, addition FP32 FP16 BFloat16 2.4 (1.7) $\textrm{mm}^{2}$ 1.1 (0.8) $\textrm{mm}^{2}$ 1.0 (0.8) $\textrm{mm}^{2}$ 22.9 mW 8.4 mW 7.3 mW 100 MHz
@@
We finally synthesized the IterL2Norm macro for each of FP32/FP16/BFloat16 using the Synopsys SAED 32/28nm technology PDK with a supply voltage of 1.05V and a clock frequency $f_{\text{clk}}$ of 100MHz. We used the Design Compiler V-2023.12-SP5. Note that the configuration of Input buffer banks for all formats follows the generic architecture (use of eight Input buffer banks, each of which stores $16\times 8$ input elements) explained in Sec. IV. The synthesis results are summarized in Table II. As such, the macro for FP32 needs on-chip memory (96.5 kib) twice as large as those (48.3 kib) for FP16 and BFloat16. For FP32, each of the Input, $\gamma$, and $\beta$ buffers uses 32 kib to store 1024 elements maximally, and the partial sum buffer uses 0.5 kib to maximally store 16 partial sums. For FP16 and BFloat16, the memory usage is half that for FP32 such that the Input, $\gamma$, and $\beta$ buffers use 48 kib in total, and the partial sum buffer 0.25 kib.
@@
As shown in Table II, the macro areas for FP32, FP16, and BFloat16 are 2.4, 1.1, and 1.0 $\textrm{mm}^{2}$, respectively. The area breakdown for each format is shown in Figs.6a-c. For all formats, the memory (Input/$\gamma$/$\beta$ and partial sum buffers) occupies the largest area in the macro, which is followed by the logic area including FP multipliers and adders. Although we considered FP multipliers and adders dedicated solely to IterL2Norm, IterL2Norm can use them in the MatMul block co-integrated on the same die. Therefore, the actual area of the IterL2Norm macro likely excludes the multiplier and adder areas, which is also listed in Table II. The operational power is also primarily determined by the FP multipliers and adders, resulting in the highest (lowest) power consumption for FP32 (BFloat16) as identified in Table II and power breakdown for each format in Figs.6d-f.